# Dredge Word, Social Media, and Webgraph Networks for Unreliable Website Classification and Identification

### v1

@@ 4.1 Single Mode Graphs
In the $\mathcal{G}_{w}$, the constructed webgraph, we define $Z$ as the domain-level attributes extracted from Ahrefs. Following Carragher et al. (2024), we log-normalize the features of each domain. However, the social media network, $\mathcal{G}_{t}$ presents a challenge in defining $Z$, as users do not have website-level attributes. For this graph, we therefore define $Z_{n}$ as features calculated using using Node2Vec Grover and Leskovec (2016) parameterized with a walk length to 20, context size to 10, and 10 walks per node, and an embedding dimension of 23, chosen because it is identical to the number of features that we extract for each website from Ahrefs. $\mathcal{G}_{t}$ presents an additional challenge in that not all websites $\in\mathcal{W}$ are mentioned by users $\in\mathcal{U}$. We include social-media-only baselines, but these are trained and evaluated on only a network of user interactions with the 2,475 labeled domains mentioned in the social media data, which make these results not directly comparable with our other models.
4.2 Homogeneous Graph
We construct a homogeneous graph $\mathcal{G}_{Hom}=\mathcal{G}_{w}\cup\mathcal{G}_{t}$ where $\cup$ denotes a graph union operation. As node features for users and websites are different, we consider only features derived from the graph structure in the homogeneous model. Again, as features are different for users and websites, we calculate features using node2vec over the graph $\mathcal{G}_{Hom}$ using the same parameters as for the single-mode graph.
4.3 Heterogeneous Graph Construction
@@
We construct a heterogeneous graph $\mathcal{G}_{Het}=(\mathcal{W}\cup\mathcal{U},A,Z,\phi,Y)$ where $\phi$ denotes a type of relation: websites-to-websites or users-to-websites. As a previous work has demonstrated the effectiveness of webgraph features in reliability classification Carragher et al. (2024), we consider only two different feature combinations. We consider a heterogeneous graph $\mathcal{G}_{Het}$ where websites have log-normed attributes from ahrefs and users are assigned the node2vec attributes calculated for $\mathcal{G}_{Hom}$. We consider a second heteroegeneous graph, $\mathcal{G}_{Het,u}$ where each user is assigned attributes extracted from their text content. For each user, we randomly sample, without replacement, 10 tweets for each observed user and embed them using multilingual distilBERT Sanh et al. (2019) [^6] All models in this paper were run or trained on a single NVIDIA GeForce RTX 3080.. This is a naive approach that excludes a massive amount of context for some users, particularly as we observed two users tweet over 1 million times in our filtered dataset. Clearly, only considering 10 tweets is suboptimal, but it is a standard practice in user embedding literature Pan and Ding (2019). A clear path for future work is to explore the impact more advanced user embedding strategies within this system.
4.4 Heterogeneous Dredge-word Graph Construction
We construct an additional heterogeneous graph $\mathcal{G}_{Het-d}$ that attempts to model the path through which dredge words can drive users to unreliable content on search engines. $\phi$ in this graph can denote connections between dredge words and websites extracted from SERPs or connections between dredge words and users $\in\mathcal{U}$. Dredge words embeddings were extracted using multilingual distilBERT, and these embeddings are used as features for dredge word nodes. Though the dredge word Twitter dataset was collected separately, but we choose to restrict the users we consider to those who also shared links to at least one known unreliable domain to facilitate model comparability. We finally define $\mathcal{G}_{Het-Dredge}$ as $\mathcal{G}_{Het-d}\cup\mathcal{G}_{Het}$.
4.5 Graph Neural Networks
@@
A label-stratified 80/10/10 split on labeled websites is used to create training, validation, and test sets. For our one-mode, heterogeneous, and homogeneous experiments we use a simple 2-layer graph neural network using GraphSAGE convolutions proposed in Hamilton et al. (2017). We ran experiments using graph attention networks Veličković et al. (2017) and heterogeneous graph transformers Hu et al. (2020), but found marginal accuracy gains at the cost of longer training times. Consequently, we selected GraphSAGE layers. Each model consists of a SAGEConv layer with dropout, with a hidden dimension of 512, followed by a ReLU activation and a second SAGEConv layer with a log softmax activation function. For each model, we train for a maximum of 1,000 epochs with early stopping based on validation loss and a patience of 50. We use an Adam optimizer with a starting learning rate of 1e-3 and a cosine annealing learning rate scheduler. The heterogeneous GNN, the most complex network we consider, contains 3.2M parameters, and no model exceeded 5 minutes of training time to reach our early stopping convergence condition.
4.6 Curriculum Learning
We assume that highly reliable and highly unreliable websites are easier to differentiate in webgraphs than those that are mixed, mostly reliable, or mostly unreliable. A manual examination of the labeled URL data leads us to believe that this assumption is very reasonable when website content is considered (see appendix B).
@@ 4.6 Curriculum Learning
We implement a slightly-modified version of the "Baby Steps" learning curriculum proposed and explored in Spitkovsky et al. (2010); Cirik et al. (2016). As domain labels from Lin et al. (2023) contain unified principal component scores of expert ratings of domain reliability, we can develop a curriculum that first learns labels of extremely reliable and extremely unreliable domains, and that gradually works from the extremes towards the reliability boundary. Using the original labeled domain dataset $\mathcal{D}$, we calculate quartiles of reliable labels and unreliable labels using principal component scores, and define these as an ordered set of batches $\{d_{1},d_{2},\dots,d_{10}\}$, where $d_{1}$ is the first quintile of the most reliable domains and $d_{10}$ is the fifth quintile of the unreliable domains. When using curriculum learning, we begin training the model $\mathcal{M}$ using $\{d_{1},d_{10}\}$, and following convergence, the model considers $\{d_{1},d_{10}\}\cup\{d_{2},d_{9}\}$. The model continues in this fashion until it converges on all available data. In our Baby Steps implementation, the we define model’s corresponding convergence as 10 epochs without an improvement in validation loss. We note that this is the curriculum convergence condition, and not the convergence condition of the model. More generally, for a website reliability curriculum $\mathcal{C}$, an even-length dataset $\mathcal{D}$, and a model $\mathcal{M}$, our implementation of the Baby Steps curriculum can be expressed as the procedure in Algorithm 1.
@@
      $B\leftarrow B\cup\mathcal{D}[i]\cup\mathcal{D}[\frac{k}{2}-i]$
@@
4.7 Unreliable Domain Discovery
@@
We compare our discovery processes with the webgraph-based discovery (WG-BD) process proposed in Carragher et al. (2024) and with the social media-based discovery process (SM-BD) proposed in Chen and Freire (2020). Specifically, we consider Precision@5, @10, @20, and we consider the partial F1 metric proposed in Chen and Freire (2020). This means we run the discovery process twice, once on the full domain list where we evaluate results manually with top-10 and top-20 accuracies, and again on a restricted domain list to compute the partial F1 metric. We additionally report partial precision and recall. We note that dredge words were only considered for 46 of the least reliable websites, so their inclusion in models is likely to harm the models, as those features only exist for a small set of the 11,327 labeled domains. We nonetheless report statistics and discovery evaluations.
Due to a lack of ground truth labels for evaluating newly discovered domains, Partial F1 measures the ability of the discovery system to find unreliable domains with respect to two known lists of unreliable sources, a seed list and an evaluation list. In this case, the seed list is the PoliticalNews dataset and the evaluation list is drawn from unreliable MBFC domains (Chen and Freire, 2020). As such, partial F1 measures the ability of a discovery system seeded on PoliticalNews to discover as many MBFC domains as possible (high partial recall), without discovering domains that are not in the MBFC list (high partial precision) (Castelo et al., 2019). We discuss the limitations of the partial F1 metric in more detail in Limitations and Appendix C.
5 Results
5.1 Credibility Classification
We present Accuracy and F1 statistics for GNN ablations in Table 1. $\mathcal{H}$ denotes a homogeneous GNN and $\mathcal{E}$ denotes a heterogeneous GNN. The subscripts $u$ and $w$ denote the use of the user network ($u$) and/or the webgraph network ($w$). We use graph-specific ($\mathcal{G}_{w}$, $\mathcal{G}_{t}$, or $\mathcal{G}_{w}\cup\mathcal{G}_{t}$) Node2Vec features in all GNNs without $f$ or $v$ subscripts. $f$ denotes the inclusion of webgraph features, $v$ denotes the inclusion of embedded user text, and $d$ denotes the model was trained on $\mathcal{G}_{Het-Dredge}$. Finally, $c$ denotes the inclusion learning curriculum in training. We note that the two social-media-only models, $\mathcal{H}_{u}$ and $\mathcal{H}_{u,c}$ are trained and evaluated on only the 2,475 labeled domains that were mentioned in the cleaned social media data.
We find that the heterogeneous model $\mathcal{E}_{u,w,c}$, which incorporates the user network, the webgraph, and curriculum learning outperforms all other models with an F1 of 0.7777±.003plus-or-minus0.7777.0030.$77$003. While the accuracy of the Social Media models is higher, those models are run only on a subset of the data, and thus are not directly comparable.
@@ 5.1 Credibility Classification
$\mathcal{H}_{w}$ 0.7686 ± .004 0.7545 ± .005
$\mathcal{H}_{w,f}$ 0.7799 ± .002 0.7714 ± .002
$\mathcal{H}_{w,f,c}$ 0.7791 ± .002 0.7705 ± .002
$\mathcal{H}_{u}^{*}$ 0.8012* ± .005 0.7443* ± .008
$\mathcal{H}_{u,c}^{*}$ 0.8113* ± .007 0.7582* ± .009
$\mathcal{H}_{u,w}$ 0.7523 ± .005 0.7654 ± .005
$\mathcal{H}_{u,w,c}$ 0.7696 ± .006 0.7559 ± .006
$\mathcal{E}_{u,w}$ 0.7797 ± .010 0.7707 ± .012
$\mathcal{E}_{u,w,c}$ 0.7865 ± .002 0.$7$003
$\mathcal{E}_{u,w,f,v}$ 0.7827 ± .007 0.7741 ± .008
$\mathcal{E}_{u,w,c,f,v}$ 0.7738 ± .118 0.7657 ± .018
$\mathcal{E}_{u,w,c,f,v,d}$ 0.7787 ± .004 0.7675 ± .007
Table 1: Mean Accuracy and F1 and standard deviations for each GNN ablation over 10 runs. $\mathcal{H}$ denotes homogeneous and $\mathcal{E}$ denotes a heterogeneous and subscripts denote included networks and features. * denotes social-media-only—these were run and evaluated on the 2,475 labeled domains mentioned in the Twitter data and thus are not directly comparable with other models.
5.2 GNN Discovery
Two annotators77725-30 year-old PhD candidates studying misinformation independently annotated the reliability of the top 20 predictions (sorted by prediction confidence) of the top performing model, $\mathcal{E}_{u,w,c}$, over the set of all unlabeled domains. Inter-annotator agreement in the first round had a Krippendorff’s $\alpha=0.78$. The annotators then met and resolved all disagreements. We compare our model with the webgraph-based discovery (WG-BD) approach in Carragher et al. (2024). As that paper’s discovery process did not rank discovered domains, we sort results by the misinformation classifier proposed in the work. Again, two independent annotators ranked the results, yielding a Krippendorff’s $\alpha=0.69$. For comparison with previous unreliable domain discovery approaches, we additionally report partial F1, defined in Chen and Freire (2020) as the set of true positive unreliable domains with credibility labeled “mixed” or worse by Media Bias Fact Check. Thresholding at a prediction confidence level of 0.7 (see Appendix C for sensitivity analysis), the Partial F1 of our top-performing GNN model is 0.25, which is largely in-line with Chen and Freire (2020) (0.29) and Carragher et al. (2024) (0.28). We observe that $\mathcal{E}_{u,w,c}$ outperforms both competing systems at each level of precision we consider.
5.3 Dredge Word Discovery
@@
[…]
6 Conclusion
We propose a model for unreliable domain detection that incorporates many of the diverse paths that users take to unreliable domains. We propose the idea of Dredge Words and highlight their bidirectional connections with social media. Finally, we demonstrate that our best-performing proposed model outperforms competing systems on the task of unreliable domain discovery.

### v2

@@ Methods
As the features of domains and users are drawn from different spaces and have different sizes, we elect to use positional node features calculated individually on each of the three constructed homogeneous graphs. For this graph, we therefore define $Z_{n}$ as features calculated using using Node2Vec (Grover and Leskovec 2016) parameterized with a walk length to 20, context size to 10, and 10 walks per node, and an embedding dimension of 23, chosen because it is identical to the number of features that we extract for each website from Ahrefs.
$\mathcal{H}_{users}$ presents an additional challenge in that not all websites $\in\mathcal{W}_{target}$ are mentioned by users $\in\mathcal{U}$. We include social-media-only baselines, but these are trained and evaluated on only a network of user interactions with the 2,475 labeled domains mentioned in the social media data, which make these results not directly comparable with our other models.
Additional Baselines
We include comparisons to two past works that explored website credibility classification without including article or website content. First, we re-implemented the GNN for news domain credibility classification and discovery proposed in (Carragher, Williams, and Carley 2024a), which is a homogeneous model which uses the Ahrefs features of domains. Following the authors, we log-normalize the Ahrefs features of each domain. We train and evaluate their model on all labeled 11,327 reliability-labeled domains.
Second, we compare our results to those from (Yang and Menczer 2023), where the authors asked ChatGPT to return the probability that a website is unreliable. Following the authors recommendation, we take a binary cut-off at 0.5 and compare results accordingly. As there is no hold-out set for the ChatGPT data, we consider two comparisons between our best-performing model and this approach. We first compare accuracy and F1 predictions for all 7,318 sites; however, this comparison is problematic as this guarantees leakage—i.e., some of these sites were in our model’s training data. To account for leakage, we perform a second evaluation where we only evaluate the predictions of our model and of ChatGPT on the hold-out test set of domains that our model never encountered.
Heterogeneous Graphs
We construct two heterogeneous graphs. $\mathcal{E}_{domains+users}$ is structurally equivalent to its homogeneous counterparts, but explicitly incorporates the two node types ($\mathcal{W},\mathcal{U}$) and their different relations ($\phi_{1},\phi_{2}$). we use a heterogeneous graph-neural network architecture that can treat user nodes and domain nodes as different nodetypes. We additionally construct $\mathcal{E}_{domains+users+dredge}$, which contains all node types ($\mathcal{W},\mathcal{U},\mathcal{Q}$) and all relation types ($\phi_{1:4}$).
@@
In the heterogeneous networks, the features of domain nodes $\mathcal{W}$ are always the logged domain-level features extracted from Ahrefs, as in Carragher, Williams, and Carley (2024a). For user nodes, we ran each model with positional context—Node2Vec features, and separately with text context. To create textual context, for each user, we randomly sample, without replacement, 10 tweets for each observed user and embed them using multilingual distilBERT (Sanh et al. 2019). This is a naive approach that excludes a large amount of context for some users, particularly as we observed two users tweet over 1 million times in our filtered dataset. However, considering only 10 tweets is a standard practice in user embedding literature (Pan and Ding 2019). A clear path for future work is to explore the impact of more advanced user embedding strategies within this system. Dredge word embeddings were extracted using multilingual distilBERT, and these embeddings are used as features for dredge word nodes. We note that dredge words were only considered for 46 of the least reliable websites, so their inclusion in models is likely to harm the models, as those features only exist for a small set of the 11,327 labeled domains. We nonetheless report statistics and discovery evaluations.
Graph Neural Network Training
@@
A label-stratified 80/10/10 split on labeled websites is used to create training, validation, and test sets. Nodes that did not have reliability labels—the majority of nodes in all networks—were masked during training. For our one-mode, heterogeneous, and homogeneous experiments we use a simple 2-layer graph neural network using GraphSAGE convolutions proposed in (Hamilton, Ying, and Leskovec 2017). We ran experiments using graph attention networks (Veličković et al. 2017) and heterogeneous graph transformers (Hu et al. 2020), but found marginal accuracy gains at the cost of longer training times. As our primary classification interest is evaluating the impact of context—i.e., signals from different sources—on domain credibility classification, rather than finding the best architecture, we elect to use homogeneous and heterogeneous GraphSAGE layers for each respective model. We include a single baseline that uses a heterogeneous graph transformer as well. Each model consists of a SAGEConv layer with dropout, with a hidden dimension of 512, followed by a ReLU activation and a second SAGEConv layer with a log softmax activation function. For models trained with user text, we included an additional linear layer to align input feature sizes. For each model, we train for a maximum of 1,000 epochs with early stopping based on validation loss and a patience of 50. We use an Adam optimizer with a starting learning rate of 1e-3 and a cosine annealing learning rate scheduler [^7] Implementation is available at ¡removed for anonymity¿. The heterogeneous GNN, the most complex network we consider, contains 3.2M parameters, and no model exceeded 5 minutes of training time to reach our early stopping convergence condition. All models in this paper were run or trained on a single NVIDIA GeForce RTX 3080.
Curriculum Learning
@@
We assume that highly reliable and highly unreliable websites are easier to differentiate in webgraphs than those that are mixed, mostly reliable, or mostly unreliable. A manual examination of the labeled URL data leads us to believe that this assumption is very reasonable when website content is considered (see Appendix B). We implement a slightly-modified version of the“Baby Steps” learning curriculum proposed and explored in Spitkovsky, Alshawi, and Jurafsky (2010); Cirik, Hovy, and Morency (2016). As domain labels from Lin et al. (2023) contain unified principal component scores of expert ratings of domain reliability, we can develop a curriculum that first learns labels of extremely reliable and extremely unreliable domains, and that gradually works from the extremes towards the websites with mixed reliability. Using the original labeled domain dataset $\mathcal{D}$, we calculate quintiles of reliable labels and unreliable labels using principal component scores, and define these as an ordered set of batches $\{d_{1},d_{2},\dots,d_{10}\}$, where $d_{1}$ is the first quintile of the most reliable domains and $d_{10}$ is the fifth quintile of the unreliable domains. When using curriculum learning, we begin training the model $\mathcal{M}$ using $\{d_{1},d_{10}\}$, and following convergence, the model considers $\{d_{1},d_{10}\}\cup\{d_{2},d_{9}\}$. The model continues in this fashion until it converges on all available data. In our Baby Steps implementation, the we define model’s corresponding convergence as 10 epochs without an improvement in validation loss. Once the curriculum has incorporated all data, the convergence criterion described in the previous section are used. More generally, for a website reliability curriculum $\mathcal{C}$, an even-length dataset $\mathcal{D}$, and a model $\mathcal{M}$, our implementation of the Baby Steps curriculum can be expressed as the procedure in Algorithm 1.
@@
      $B\leftarrow B\cup\mathcal{D}[i]\cup\mathcal{D}[k-i]$
@@
Unreliable Domain Discovery
While we can evaluate our classifiers on domains in $\mathcal{W}_{target}$, we desire to make a tool that can identify unknown unreliable domains. To evaluate our best-performing classifier, we explore its ability to identify unlabeled unreliable domains over the unlabeled domains in $\mathcal{W}_{source}$. We outline our discovery processes and we benchmark our discovery process against two previous works.
@@
We compare our discovery processes with the webgraph-based discovery (WG-BD) process proposed in Carragher, Williams, and Carley (2024a) and with the social media-based discovery process (SM-BD) proposed in Chen and Freire (2020). Specifically, we consider Precision@5, @10, @20, and we consider the partial F1 metric proposed in Chen and Freire (2020). This means we run the discovery process twice, once on the full domain list where we evaluate results manually with top-10 and top-20 accuracies, and again on a restricted domain list to compute the partial F1 metric. We additionally report partial precision and recall.
Due to a lack of ground truth labels for evaluating newly discovered domains, Partial F1 measures the ability of the discovery system to find unreliable domains with respect to two known lists of unreliable sources, a seed list and an evaluation list. In this case, the seed list is the PoliticalNews dataset and the evaluation list is drawn from unreliable MBFC domains (Chen and Freire 2020). As such, partial F1 measures the ability of a discovery system seeded on PoliticalNews to discover as many MBFC domains as possible (high partial recall), without discovering domains that are not in the MBFC list (high partial precision) (Castelo et al. 2019). Following Carragher, Williams, and Carley (2024a), we discuss the limitations of the partial F1 metric in more detail in Limitations and Appendix C.
Results
Credibility Classification
We present Accuracy and F1 statistics for website reliability classification for homogeneous models ($\mathcal{H}$) and heterogeneous models ($\mathcal{E}$) in Table 1. We find that the heterogeneous model which incorporates both the user network and domain webgraph outperforms all other models with an average accuracy over 10 runs of $0.7865\pm.002$. However, swapping the Heterogeneous GraphSage Convolutions in this model with a Heterogeneous Graph Transformer ($HetGT$) convolutions yielded a slightly better F1 ($0.789\pm.003$). While the accuracy of the Social Media user model ($\mathcal{H}_{users}$ is higher, this model is only evaluated the 2,475 labeled domains in $\mathcal{W}_{target}$ to which social media users were connected, and thus are not directly comparable to other models. In the heterogeneous models, we considered embedded user text features (text) as well as user node node2vec features (n2v). Interestingly, including embedded text of social media users as node performed worse than using features that uniquely consider network position.
@@ Results
$\mathcal{H}_{domains}$ 0.7686 ± .004 0.7545 ± .005
Carragher et al. 0.7799 ± .002 0.7714 ± .002
$\mathcal{H}_{users}^{*}$ 0.8113* ± .007 0.7582* ± .009
$\mathcal{H}_{domains+users}$ 0.7696 ± .006 0.7559 ± .006
$\mathcal{E}_{domains+users(n2v)}$ 0.7865 ± .002 0.$7$003
$HetGT_{domains+users(n2v)}$ 0.7820 ± .003 0.7895 ± .003
$\mathcal{E}_{domains+users(text)}$ 0.7738 ± .118 0.7657 ± .018
$\mathcal{E}_{domains+users+dredge}$ 0.7787 ± .004 0.7675 ± .007
Table 1: Mean Accuracy, F1, and standard deviations for each GNN ablation over 10 runs. $\mathcal{H}$ denotes homogeneous and $\mathcal{E}$. * denotes a social-media-only model run and evaluated on the 2,475 labeled domains mentioned in the Twitter data.
We compare our best performing model with the approach proposed in Yang and Menczer (2023) in Table 2. The authors in Yang and Menczer (2023) used ChatGPT to return the probability that 7,318 websites—7,317 of which are in our list—were unreliable labeled domains. Following the authors, we use a cutoff of 0.5 to binarize their ChatGPT predictions. Our model was trained on many of the 7,318 websites that GPT evaluated, so (full data) comparisons result in label leakage. Consequently, we provide a second (test set) evaluation, where we only compare labeled websites that our models never encountered in training. On both the full dataset and the test set evaluations, our model outperformed ChatGPT. Our model yielded an F1 that was 0.03 higher than ChatGPT on the full dataset and 0.04 higher on the test set.
Accuracy F1
GPT (full data) 0.81 0.751
$\mathcal{E}_{domains+users(n2v)}$ (full data) 0.837 0.78
GPT (test set) 0.782 0.701
$\mathcal{E}_{domains+users(n2v)}$ (test set) 0.819 0.745
Table 2: Our best model outperforms the LLM approach proposed in Yang and Menczer (2023). We compare against all 7,318 websites the authors considered (full data) as well as only the websites our model never encountered during training (test set).
Unreliable Domain GNN Discovery
Two annotators88825-30 year-old PhD candidates studying misinformation independently annotated the reliability of the top 20 predictions (sorted by prediction confidence) of the top performing model, heterogeneous model domains+users(n2v), over the set of all unlabeled domains. Inter-annotator agreement in the first round had a Krippendorff’s $\alpha=0.78$. The annotators then met and resolved all disagreements. We compare our model with the webgraph-based discovery (WG-BD) approach in Carragher, Williams, and Carley (2024a) [^9] As that paper’s discovery process did not rank discovered domains, we sort results by the misinformation classifier proposed in the work.. Again, two independent annotators ranked the results, yielding a Krippendorff’s $\alpha=0.69$.
For comparison with previous unreliable domain discovery approaches, we additionally report partial F1, defined in Chen and Freire (2020) as the set of true positive unreliable domains with credibility labeled “mixed” or worse by Media Bias Fact Check. Thresholding at a prediction confidence level of 0.7 (see Appendix C for sensitivity analysis), the Partial F1 of our top-performing GNN model is 0.25, which is largely previous works; 0.29 for SM-BD (Chen and Freire 2020)) and 0.28 for WG-BD (Carragher, Williams, and Carley 2024a). As demonstrated in Table 3, we observe that our heterogeneous domains+users model outperforms both competing systems at each level of precision we consider.
Dredge Word Discovery
@@
In the latter approach, as a consequence of 1) only extracting dredge word networks for only 46 of the least reliable domains 2) the dominance of social media and shopping sites in the dredge word SERPs, the dredge word GNN learns to heavily associate social media sites and online shopping platforms with unreliable domains. The 10 most confident predictions the GNN returns contain 7 social media sites, amazon, itunes, and s3.amazonaws. We find at least several of the pseudo-scientific and conspiratorial dredge words are targeted by dubious sellers, podcasters, and social media influencers. As these websites are out-of-domain for our task, but still often contain unreliable content, we chose not to evaluate Precision@k for this set. A deeper exploration of the link between these conspiratorial dredge words and non-news domains would be a fruitful path for future work.
@@
Table 3: Precision@K and Partial F1 (PF1), Partial Precision (PP), and Partial Recall (PR) discovery evaluations. (Chen and Freire 2020).
Analysis and Discussion
We explore the 25 most confident unreliable predictions of our best performing model over the set of all unlabeled domains in $\mathcal{W}_{source}$. We observe that these predictions seem to fall into three broad categories: health misinformation (6 websites), Qanon [^10] A conspiracy theory that claims that a cabal of Satanist, sex-trafficking, child molesters is running the world. misinformation (5), and German-language far-right misinformation (9). In all three categories, the majority of the websites express skepticism towards vaccines, the existence of COVID-19, or both. Most interestingly, our investigation of the these categories captures an interplay between websites spreading medical misinformation and the websites profiting off of it.
Six of our model’s 25 most confident predictions, are an likely co-owned network of websites selling dubious medical products. The $\mathcal{W}_{target}$ website “holitichealth” [^11] https://web.archive.org/web/20240329171920/https://holistichealth.one/ offers users advice on how to detox from vaccines, which the site claims “poison your DNA, brain, nervous system and immune system”. The six sites that most frequently link to holistichealth all appeared in our model’s most confident unreliable predictions—all sites have the same HTML template, end in .one, and resolve to the same IP address, and frequently link to one another, heavily suggesting co-ownership. Our model’s second most confident prediction, herbalremedies [^12] https://web.archive.org/web/20240225051441/https://herbalremedies.one/herbs-for-cancer/, sells 17 supplements that the site claims kill cancer cells, including colloidal silver and ”parasite cleansing herbs”. Another .one site, “emfprotection”, sells electric and magnetic field protection gear that the site claims will defend customers from psychic mind control and 5G radio waves. While we did not observe any Twitter mentions of the 6 unlabeled sites, holitichealth, was mentioned by 7 Twitter accounts. Interestingly, the 7 Twitter accounts often linked to some reliable news sites, but also linked to Zero Hedge, The Epoch Times, Breitbart, and other low-and-mixed reliability websites.
The second large category of discovered sites promoted conspiratorial—often Qanon (Garry et al. 2021)—content. Three of these sites (wakeup.icu, vintel1776.net, and qaggregator.news), are now dead, but these sites were all active during the period that Twitter data were collected. Two of the sites linked heavily to a website containing all of Q’s drops [^13] https://web.archive.org/web/20240911060604/https://qagg.news/, and wakeup linked heavily to a Qanon discussion board. The only identified Qanon website that was mentioned on Twitter was Hnewswire [^14] https://web.archive.org/web/20240903154602/https://hnewswire.com/, which was mentioned by 100 distinct Twitter accounts that also heavily mentioned Zero Hedge, Gateway Pundit, and many other low and mixed reliability domains. Hnewswire’s navigation menu contains a trending topics bar which includes categories like “Covid Kill Shot”, “Plandemic”, and “Demonic Activity End Times”. Independent annotations of predictions can be found in the Github repository.
The top 25 discovered websites were mentioned by 173 Twitter users in the COVID data who cumulatively mentioned over 3,165 domains. Interestingly, 42 of these users also used dredge words at least once in COVID-19-related tweets during that time period. Most of these are names of unreliable websites, e.g., “creation”, “gateway pundit”, “rense”, “infowars”, and “bitchute”, but the dredge words “what really happened”, “national vaccine”, “being 6”, and “missing links” were also used by this set of users.
This analysis highlights the connectivity of COVID-19 misinformation. Both the Qanon websites and the far-right German-language websites (which we do not discuss due to space constraints) actively promoted COVID misinformation, often alongside other conspiracies. This misinformation was then monetized by the .one websites, which sell, for example, “natural Hydroxychloroquine and Ivermectin alternative[s]” to those concerned about COVID [^15] https://web.archive.org/web/20240420185505/https://holistichealth.one/natural-hydroxychloroquine-and-ivermectin/. This is a key advantage of the current method over article-based approaches; incentive structures are often built into networks. There is often overlap in those spreading conspiracies and those monetizing them (Ballard et al. 2022), and having a system that considers the diverse paths that users can take to unreliable content helps us identify that interplay.
Limitations
While we attempted to mitigate limitations where possible, several substantial ones exist. First, the Twitter data were collected from 2020-2022, but the dredge words used in this paper were based on Google SERPs in 2024. We cannot say that the dredge words we extracted surfaced the same content in 2024 as when the phrases were initially used on Twitter. As with the Ukraine Biolabs example, there are also likely many cases where unreliable websites ranked highly, but were overtaken by fact checkers; our approach would not capture this. Future work could explore the impacts of better temporal alignment, which would have the added benefit of covering data voids formed around breaking news stories.
A key limitation is that the current process for extracting dredge words relies on paid third-party services. While we validated each of the extracted dredge words to ensure they are ranking in Google’s top-10, approaches to extract keyphrases from, for example, common-crawl webgraphs would provide more democratization of this line of research. We publish all dredge word data collected for this project in line with that goal. Additionally, the Partial F1 metric we use for discovery evaluation is limited, as previous work has shown many PoliticalNews and MBFC sites on which the metric relies are dead or no longer active (Carragher, Williams, and Carley 2024a). However, to our knowledge, no better metric exists for evaluating unreliable domain discovery systems. This is a needed avenue for future work.
Conclusion
Proactive content moderation requires systems that rapidly and continuously identify unreliable information sources. By considering additional direct paths that users can take to unreliable websites, we can improve the ability of models to identify these sources. We propose the idea of Dredge Words and highlight their bidirectional connections with social media. Finally, we demonstrate that our best-performing model outperforms competing systems on the tasks of domain reliability classification and unreliable domain discovery. This work demonstrates the signals present in the direct and indirect pathways that users follow to unreliable websites. Better understanding these paths could be a promising area of collaboration between platforms, organizations, and researchers intent on mitigating the spread of misinformation across digital ecosystems.